A 20-kg wagon is pulled along the level ground by a rope inclined at 30° above the horizontal. A friction force of 30 N opposes the motion. How large is the pulling force if the wagon is moving with an acceleration of 0,40 m/s2? Give the answer with one decimal and no units.
1
Expert's answer
2020-10-30T13:12:11-0400

The horizontal component of the force is:


"F_h = F\cos 30\degree = \dfrac{F\sqrt{3}}{2}"

where "F" is the required pulling force. According to the second Newton's law along the horizontal line, have:


"F_h-F_{fr} = ma"

where "F_{fr} = 30N" is the friction force, "m = 20kg" is the mass and "a = 0.4m\/s^2" is the accleration of the wagon. Substituting the expression for "F_h" and expressing "F", obtain:


"\dfrac{F\sqrt{3}}{2} - F_{fr} = ma\\
\dfrac{F\sqrt{3}}{2} = ma+F_{fr} \\
F = \dfrac{2}{\sqrt3}(ma+F_{fr})\\
F = \dfrac{2}{\sqrt3}(20\cdot 0.4+30) \approx 43.9N"

Answer. 43.9.